OpenStudy (anonymous):

Divide. (10a4 – 5a3) ÷ 5a A. 10a^4 – 5a^3 – 5a B. 10a^4 – 5a^3 + 5a C. 2a^3 – 5a^3 D. 2a^3 – a^2

jimthompson5910 (jim_thompson5910):

hint: factor out 5a from the numerator

OpenStudy (anonymous):

I got D..But I'm still confused, I don't think i got it right :/

jimthompson5910 (jim_thompson5910):

Another way you can do it is to divide each term in the numerator by 5a in the denominator First term: \[\Large \frac{10a^4}{5a} = \frac{10}{5}\frac{a^4}{a^1} = 2a^{4-1} = 2a^3\] Second term: \[\Large \frac{-5a^3}{5a} = -\frac{5}{5}\frac{a^3}{a^1} = -1a^{3-1} = -a^2\]

jimthompson5910 (jim_thompson5910):

So \[\Large 2a^3-a^2\] is correct

OpenStudy (anonymous):

Really? :D Oh wow I thought I was wrong

jimthompson5910 (jim_thompson5910):

did you follow those steps or did you take another route?

OpenStudy (anonymous):

I did it in a way that I don't even understand I would explain it to you but I don't know how, I did half of it in my head and half of it on paper because I started over twice.

jimthompson5910 (jim_thompson5910):

alright I'm just curious. Hopefully my method makes sense?

OpenStudy (anonymous):

Yes it does :) I'm gonna try that next time

jimthompson5910 (jim_thompson5910):

alright great

OpenStudy (anonymous):

Thank you so much!

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

Can I ask you one more question?

jimthompson5910 (jim_thompson5910):

sure

jimthompson5910 (jim_thompson5910):

go for it

OpenStudy (anonymous):

Determine whether the polynomial is a perfect square and if it is, factor it. n^2 – 6n + 9 A. is a perfect square: (n – 3)^2 B. is not a perfect square C. is a perfect square: (n – 6)^2 D. is a perfect square: (n + 3)^2 Would I do the same thing as the other problem?

jimthompson5910 (jim_thompson5910):

how do you determine if a polynomial is a perfect square?

OpenStudy (anonymous):

I think the first and the last term have to be squares, I'm not sure..

jimthompson5910 (jim_thompson5910):

yes that's on the right track

jimthompson5910 (jim_thompson5910):

is n^2 a perfect square?

OpenStudy (anonymous):

um I think so..

jimthompson5910 (jim_thompson5910):

it is because all perfect squares are of the form x^2 (which literally means "x squared")

jimthompson5910 (jim_thompson5910):

how about 9, is that a perfect square?

OpenStudy (anonymous):

no

jimthompson5910 (jim_thompson5910):

why not?

OpenStudy (anonymous):

Because it isn't a square root?

jimthompson5910 (jim_thompson5910):

but since 3^2 = 9, this means 9 is actually a perfect square

jimthompson5910 (jim_thompson5910):

put another way, \[\Large \sqrt{9} = 3\]

OpenStudy (anonymous):

ohhh ok, I thought it had to look something like this x^9 to be a perfect square, but now I get what you mean.

jimthompson5910 (jim_thompson5910):

no the exponents are 2 because we're dealing with squared terms

jimthompson5910 (jim_thompson5910):

so we have n^2 is a perfect square, so is 9 (since 3^2 = 9)

jimthompson5910 (jim_thompson5910):

what else do we need to check?

OpenStudy (anonymous):

The 6n part?

jimthompson5910 (jim_thompson5910):

yes, we need to see if it fits a very specific form

jimthompson5910 (jim_thompson5910):

do you remember what the form is?

OpenStudy (anonymous):

no

jimthompson5910 (jim_thompson5910):

it turns out that when you expand (a-b)^2 you get a^2 - 2ab + b^2 using the FOIL rule or you can use various other methods

jimthompson5910 (jim_thompson5910):

(a-b)^2 is a perfect square trinomial and so is a^2 - 2ab + b^2

jimthompson5910 (jim_thompson5910):

notice how in a^2 - 2ab + b^2, we have a^2 and b^2, which are perfect squares themselves the middle is -2ab

jimthompson5910 (jim_thompson5910):

so the middle is formed by multiplying the terms 'a' and 'b' and doubling it (and it happens to be negative)

jimthompson5910 (jim_thompson5910):

In this case, we have 'n' and 3 (notice how I'm using the bases), which multiply to n*3 = 3n Now double this to get 2*3n = 6n And finally make it negative to get -6n

jimthompson5910 (jim_thompson5910):

the -6n fits the form so n^2 - 6n + 9 is indeed a perfect square trinomial

jimthompson5910 (jim_thompson5910):

it is in the form (a-b)^2 = a^2 - 2ab + b^2 where a = n and b = 3

OpenStudy (anonymous):

So would it be A..?

jimthompson5910 (jim_thompson5910):

that is correct

jimthompson5910 (jim_thompson5910):

here is another example to work with z^2 - 10z + 25 we have z^2, a perfect square 25 is a perfect square since 5^2 = 25 so a = z, b = 5 Plug those into (a-b)^2 = a^2 - 2ab + b^2 to get... (a-b)^2 = a^2 - 2ab + b^2 (z-5)^2 = z^2 - 2z*5 + 5^2 (z-5)^2 = z^2 - 2*5*z + 5^2 (z-5)^2 = z^2 - 10*z + 5^2 (z-5)^2 = z^2 - 10z + 25

jimthompson5910 (jim_thompson5910):

Notice how if you change that -10z to anything else, then the trinomial won't be a perfect square

OpenStudy (anonymous):

oh now I get it, But is their any possible number you could change it to? I'm just curious :)

jimthompson5910 (jim_thompson5910):

which number are you referring to change?

OpenStudy (anonymous):

The -10z I know you said that if you change it to anything else then the trinomial won't be a perfect square, but when you said that you can't change it to anything else, do you mean that if you changed it to any other number it wouldn't be a perfect square? Sorry if that's a confusing question

jimthompson5910 (jim_thompson5910):

well if you fix the outer terms to be z^2 and 25, then the middle has to be either +10z or -10z for it to be a perfect square. If it was +10z, then z^2 + 10z + 25 factors to (z+5)^2 If it was -10z, then z^2 - 10z + 25 factors to (z-5)^2

jimthompson5910 (jim_thompson5910):

however, if you change either outer term, then you can change the middle term based on what the outer terms change to

jimthompson5910 (jim_thompson5910):

for example, if you had z^2 for the first outer term and then had 36 instead of 25, then the middle term would be 2*z*6 = 12z (either plus or minus of that)

jimthompson5910 (jim_thompson5910):

meaning that z^2 - 12z + 36 = (z-6)^2 and z^2 + 12z + 36 = (z+6)^2

OpenStudy (anonymous):

ohh ok, sorry that I asked that was probably a dumb question XD But could y^2 be a perfect square?

jimthompson5910 (jim_thompson5910):

y^2 is a perfect square

jimthompson5910 (jim_thompson5910):

any variable squared is a perfect square

jimthompson5910 (jim_thompson5910):

and don't be afraid to ask questions, there's no such thing as a dumb question

OpenStudy (anonymous):

oh ok :) I'm doing a problem and it says y^2-16 and I know I'm supposed to see if the first term is a perfect square, which you said it is, and after that I/m supposed to see if the last ones a perfect square right?

jimthompson5910 (jim_thompson5910):

so is 16 a perfect square?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

so what number replaces x in the equation x^2 = 16

jimthompson5910 (jim_thompson5910):

in other words, what number do you square to get 16

OpenStudy (anonymous):

ohh ok that makes more sense, um 2?

jimthompson5910 (jim_thompson5910):

2 squared = 2^2 = 2*2 = 4

jimthompson5910 (jim_thompson5910):

so that's not it

jimthompson5910 (jim_thompson5910):

3 squared = 3^2 = 3*3 = 9

OpenStudy (anonymous):

what about 8?

jimthompson5910 (jim_thompson5910):

8^2 = 8*8 = 64

jimthompson5910 (jim_thompson5910):

see how I'm squaring any given number?

OpenStudy (anonymous):

yeah

jimthompson5910 (jim_thompson5910):

visually, "squaring" means you find the area of a square with a given side length example: find the area of a square with side length of 6